Problem #1
Student Loans Student borrowers now have more options to
choose from when selecting repayment plans.* The standard
plan repays the loan in 10 years with equal monthly payments.
The extended plan allows from 12 to 30 years to repay the loan.
A student borrows $35,000 at 7.43% compounded monthly.

Find the monthly payment and total interest paid under the
standard plan.

Problem #2
Inheritance Sandi Goldstein has inherited $25,000 from
her grandfather’s estate. She deposits the money in an
account offering 6% interest compounded annually. She
wants to make equal annual withdrawals from the account
so that the money (principal and interest) lasts exactly
8 years.
a. Find the amount of each withdrawal.
b. Find the amount of each withdrawal if the money must
last 12 years.

Answer by stanbon(48551)

The standard plan repays the loan in 10 years with equal monthly payments.
A student borrows $35,000 at 7.43% compounded monthly.
Find the monthly payment and total interest paid under the
standard plan.
-----------------
Formula:
Loan Balance = (original amount)(1+r)^n-[(payment amt)/i][(1+i)^n-1]
i is the interest for each payment period; n is # of payments
You want the loan balance to be zero.
0 = 35000(1+.0743/12)^(10*12) - [P/(0.0743/12)][(1+(0.0743/12)^120 -1]
0 = 35000(1.006191667)^120 - [P/0.006191667][1.006191667^120 - 1]
0 = 35000 - [P/0.006191667][1.0974225777}
0=73409.79 -P*177.24
P = $414.18 (monthly payment made for 120 months
Total Interest = 120*414.18-35000= $14,701.44
================================================
Inheritance Sandi Goldstein has inherited $25,000 from
her grandfather’s estate. She deposits the money in an
account offering 6% interest compounded annually. She
wants to make equal annual withdrawals from the account
so that the money (principal and interest) lasts exactly
8 years.
a. Find the amount of each withdrawal.
Same formula as above. By the way you can find formulas using Google
and searching on "investment formulas".
0 = 25000(1.06)^8-[P/0.06][(1.06)^8 - 1]
Solve for P.
----------------
b. Find the amount of each withdrawal if the money must
last 12 years.
Let n=12 and solve for P.
---------------
Cheers,
Stan H.